A current of 7.49×104 A is passed through an electrolysis cell containing molten NaCl for 9.40 days.

(a) How many grams of sodium are produced?

g



(b) How many liters of chlorine are collected, if the gas is at a temperature of 273 K and a pressure of 1.00 atm?

L
Answer #1

a.   

Na^+ (aq) + e^- ------------> Na(s)

Z = 1

M   = 23g/mole

C   = 7.49×104 A

t    = 9.4 days   = 9.4*24*60*60   = 812160sec

W     = MCt/ZF

W    = 23*7.49*10^4 *812160/(1*96500)    = 14498529g

b.

2Cl^- (aq) -------------. Cl2(g) + 2e^-

Z = 2

W    = MCt/ZF

W/M    = Ct/ZF

n        = Ct/ZF                    [ n   = W/M]

n        = 7.49*10^4 *812160/(2*96500)    = 315185moles

1 mole of any at STP condition ( 273K, 1atm) occupied volume is 22.4L

315185 moles of any STP (273K,1atm) occupied volume   = 22.4*315185/1    = 7060144L